2. Calculate the probability from the z-table. a. P(Z > -0.45) b. P(Z < 2.10) c. P(0 < Z < 2.10) d. P(-0.45 < 3 < 2.10)
Answer #1

a) P( Z > -0.45) = 1 - P( Z <= -0.45)

= 1 - 0.3264

= 0.6736

b) P( Z < 2.10 ) = 0.9821

c) P( 0 < Z < 2.10) = P( Z < 2.10) - P( Z < 0)

= 0.9821 - 0.5000

= 0.4821

d) P( -0.45 < Z < 2.10 ) = P( Z < 2.10) - P( Z < -0.45)

= 0.9821 - 0.3264

= 0.6558

(All values are obtained from normal distribution Z - table)